OpenStudy (anonymous):

Find the area of the surface obtained by rotating the curve y= \sqrt{3 x} from x=0 to x = 3 about the x-axis.

OpenStudy (anonymous):

first draw the curves

OpenStudy (anonymous):

Doin'

OpenStudy (anonymous):

How precise do I need to be?

OpenStudy (anonymous):

I have 0 and 3 when I draw the sqrt function.

OpenStudy (anonymous):

crude just so you see what you are doing :)

OpenStudy (anonymous):

Got it.

OpenStudy (anonymous):

then \[ S=\int_a^b2\pi y\sqrt{1+\left(\frac{dy}{dx}\right)^2}dx \]

OpenStudy (anonymous):

Is dy/dx the \[\sqrt{3x}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Or its derivative?

OpenStudy (anonymous):

no. y=sqrt(3x) you've gotta find dy/dx

OpenStudy (anonymous):

-3/(2*sqrt(3x))

OpenStudy (anonymous):

Okay, cool

OpenStudy (anonymous):

So, I get this gargantuan thing...

OpenStudy (anonymous):

\[\int\limits_{0}^{3} 2\pi \sqrt{3x}\sqrt{1+1/12x}\]

OpenStudy (anonymous):

Do I just plug in 3 and 0?

OpenStudy (anonymous):

the second radial has a major error

OpenStudy (anonymous):

Radial?

OpenStudy (anonymous):

What's a radial?

OpenStudy (anonymous):

\[\sqrt{1+\left(\frac{-3}{2\sqrt{3x}}\right)^2}=\sqrt{1+{3\over2x}}=\sqrt{\frac{2x+3}{2x}} \] **radical**

OpenStudy (anonymous):

\[\sqrt{3x}\sqrt{\frac{2x+3}{2x}}=\sqrt{{3\over2}(x+3)} \]

OpenStudy (anonymous):

\[ S=2\pi\sqrt{3\over2}\int_0^3\sqrt{x+3}dx \] and solve

OpenStudy (anonymous):

Where is the negative 3 coming from?

OpenStudy (anonymous):

oh yes.. my bad.. no -ve but its squared. so I got lucked out ;)

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

you are talking about the "-3" in the radical in the first step, right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

yes and it is squared.. so (-3)^2=+9 so, just ignore that -ve sign (take it as +ve like you said)

OpenStudy (anonymous):

Why is a 3 there though? Even if it's positive? 'Cause the derivative of sqrt3x is 1/2sqrt3x.

OpenStudy (anonymous):

chain rule, pal

OpenStudy (anonymous):

Chain rule? Ooooooooohhhhh....

OpenStudy (anonymous):

I'm studying for a calc final (252) and my prof. handed us 38 problems to do in two days. The little things are slipping me.

OpenStudy (anonymous):

A lot.

OpenStudy (anonymous):

gotta be careful though

OpenStudy (anonymous):

Oh yes.

OpenStudy (anonymous):

So, I get it until.....

OpenStudy (anonymous):

\[\sqrt{3x}\sqrt{2x+3/2x}\]

OpenStudy (anonymous):

I got something totally different. I got \[\sqrt{3x+9x/2x}\]

OpenStudy (anonymous):

ok that is fine too. but take the LCM of the fraction

zepdrix (zepdrix):

\[\large 2\pi \int\limits_0^3 \sqrt{3x}\sqrt{1+\left(\frac{dy}{dx}\right)^2}dx \qquad = \qquad 2\pi \int\limits_0^3 \sqrt{3x}\sqrt{1+\left(\frac{3}{2\sqrt{3x}}\right)^2}dx\] \[\large 2\pi \int\limits_0^3 \sqrt{3x}\sqrt{1+\left(\frac{9}{12x}\right)}dx \qquad = \qquad 2\pi \int\limits_0^3 \sqrt{3x}\sqrt{\frac{12x+9}{12x}}dx\] \[\large 2\pi \int\limits\limits_0^3 \sqrt{3x}\frac{\sqrt{12x+9}}{2\sqrt{3x}}dx\]

zepdrix (zepdrix):

Hmm I'm trying to figure out how you got to the spot you're at.

OpenStudy (anonymous):

i mis-wrote it.. instead of 12, i put a "2"

OpenStudy (anonymous):

Yeah I forgot the 12 too. Tell ya what...you guys reply and I'll check back later. My brain is burning out I think. I'll be back.

zepdrix (zepdrix):

\[\large \cancel2\pi \int\limits\limits\limits_0^3 \cancel{\sqrt{3x}}\frac{\sqrt{12x+9}}{\cancel2\cancel{\sqrt{3x}}}dx\qquad = \qquad \pi \int\limits_0^3 \sqrt{12x+9}dx\]

zepdrix (zepdrix):

I think you end up with something like this.

OpenStudy (anonymous):

haha. hmm .. its late

OpenStudy (anonymous):

\[ S=2\pi\int_0^3\sqrt{3x\frac{12x+9}{12x}}dx\\ =2\pi\int_0^3\frac{\sqrt{12x+9}}{2}dx\\ 12x+9=u\quad,dx=du/12\quad,9\le u\le45\\ S=\pi\int_9^{45}u^{1/2}{du\over12}={\pi\over12}{2\over3}\left[u\sqrt{u}\right]_9^{45}\\ ={\pi\over18}[45\sqrt{45}-9\cdot3]\\ S={\pi\over2}(5\sqrt{45}-3) \]